Actions
RunPlan20150721¶
July | 21 | ||
estimated time | 06:00-14:00 | 14:00-22:00 | 22:00-06:00 |
MOM | R. Bayes | R. Bayes | R. Bayes |
Lead Shifter | Jan | Celeste | Paolo |
Shifter 1 | Tanaz | Pavel | Dejan |
BLOC | Celeste | Celeste | Joe |
SOC | |||
Experimenter | |||
Aim | On Call | On Call | On Call/Tracker Commissioning |
Beam? | Yes(?) | Yes(?) | Yes(?) |
Additional Personnel | |||
Shift plans¶
Magnet Commissioning¶
- To be conducted for the majority of the day.
- Shifters will provide support for these activities in the form of
- monitoring solenoid state
- controlled entry procedures
- other activities as requested
Magnet Training¶
Two sessions today at 10:00 and 15:00 in CR09. Shifters should attend when possible.
Half Field Run¶
If a stable magnetic field is generated expect to take data- DS is required
- Propose to take data using a 6-140+M0 beam.
- Check rate at setup time
- TOF2 trigger rate should be > 5 per spill.
- Reference run is a potential fall back position if the rate is too low.
Tracker Commissioning¶
Propose to complete tracker commissioning over midnight shift if there is no field at that time.- Use Ed's script to set the tracker trigger windows for timing study.
TOF Calibration¶
If there are not Tracker timing instructions by run time and no field take TOF calibration data for the duration of the midnight shift instead.
Number of pulses | Beamloss | Particle Species | p at Tgt | p@D1 | p@D2 | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s) |
# | V | mm | A | A | A | A | A | A | A | A | A | A | A | A | A | A | A | A | DAQ run # |
25000 min | 1 | Positron | 300.0 | 259.7 | 249.58 | 29 | 56.35 | 102.7 | 62.65 | 195.72 | --- | 98.59 | 163.16 | 218.79 | 114.14 | 83.40 | 125.70 | 80.06 | 1678 |
What We Haven't Done¶
- TOF Calibration
- Have not had a block of time sufficient to accrete necessary data
- Good practice but not strictly necessary as it has not been that long since run01a.
- Tracker Timing Studies
- Ed Overton used available time for Tracker calibration.
- Has suggested producing a script to step through proposed settings for shifters to operate
Run Time Preparedness.¶
- Downstream PID Detectors (EMR/KL/TOF)
- to be moved to final positions (July 16)
- HV to be turned on and DAQ reconfigured (July 16?)
- Detectors reconnected (July 15)
- Upstream PID Detectors (Ckov/TOFs)
- in final positions -- Done (July 4)
- Trackers
- Helium on
- downstream tracker -- Done but not maintained. (July 17)
- Upstream tracker still has a leak -- to be fixed in August
- Repair dead channels.
- Helium on
- Run Control
- CDB interface
- Need to improve stability of run control - remove need for IOC restart
- CDB beam line control
- Starting beam line current from zero doesn't work due to polarimeter -- Fixed (July 12)
- CDB interface
- Beam line control
- "Set all to 0" button has not operated in a predictable manner -- Corrected. (July 12)
- Beam line
- Decay Solenoid now filled (July 18)
- To be ramped this afternoon.
- Correcting limits on the alarm handler.
- Likely to wait for September run.
- Decay Solenoid now filled (July 18)
Expected Work in the Hall¶
- Tracker commissioning and repairs to be conducted
- Will involve Paul R., Alan B. and Ed O.
- Expect them to be moving in and out of the hall
Notes¶
- Reminder: The end of run comment must be filled for the run to register as completed on CDB. Ensure that the enter key is struck while in the end comment dialogue box at the end of every run.
Temporary Run Control Instructions¶
- Restart Run Control after every run
- Currently using pro branch in MLCR.
- If Run Control was no restarted and it crashes the terminal, type "reset" (no quotes) to regain control, then restart Run Control. Do not log in/out
- If there is doubt contact Pierrick Hanlet or Chris Heidt
Phone details for 6pm daily meeting¶
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Updated by Bayes, Ryan about 8 years ago ยท 6 revisions